# Low-Dimensional Adaptation of Rectified Flow: A New Perspective through the Lens of Diffusion and Stochastic Localization

### v1

@@ 5 Experiments
We have performed extensive simulations to corroborate our theoretical findings for both simulated data and real Text-to-Image (T2I) generation experiments.
5.1 Synthetic data experiments
@@
In this section we demonstrate the usefulness of non-uniform time-grid for generation task from low-dimensional distribution. We set the intrinsic dimension $k=8$, and set our target distribution to be $p_{1}=N(\mu,\Sigma)$, where $\mu=(8,8,\ldots,8)^{\top}\in\mathbb{R}^{d}$ and $\Sigma$ is a block-diagonal matrix with blocks $I_{k}$ and $\boldsymbol{0}_{(d-k)\times(d-k)}$. In this case, the drift $v_{t_{i}}(x)$ can be computed in closed form as
$x\in\mathbb{R}^{d}\mapsto v_{t_{i}}(x)$ $=\frac{x}{t_{i}}+\frac{1-t_{i}}{t_{i}}s_{t_{i}}(x)$ (22)
$=\begin{bmatrix}\frac{2t_{i}-1}{\sigma_{t_{i}}^{2}}I_{k}&\boldsymbol{0}_{k\times(d-k)}\\ \boldsymbol{0}_{(d-k)\times k}&-\frac{1}{1-t_{i}}I_{d-k}\end{bmatrix}x.$
@@
In the experiments, we use (22) for generating new samples from $p_{1}$. We vary the ambient dimension $d$ between 10 and 800, and we let the number $N$ of discretization-steps to be $100$ and $200$. We use both a non-uniform time-grid (15) (with $\delta=\min\{1/N,1/d\}$) and a uniform time-grid for the generation task. We generate 2000 samples under the aforementioned discretization schemes and measure the TV distance with respect to another batch of 2000 samples independently generated from the slightly blurred target distribution $\text{Law}((1-\delta)X_{1}+\delta Z)$, where $Z\sim N(0,I_{d})$. We use the discriminative approach proposed in Tao et al. (2024) to estimate the TV distance between the distributions via 10 independent Monte Carlo rounds. Figure 3(a)-(b) shows that nonuniform discretization has better adaptation to the low-dimensional structure of $p_{1}$ in high-dimensional regime.
For the SDE sampler (21), we also compare its TV convergence with that of the deterministic sampler (5) with the non-uniform time grid (15). Figure 3(c)-(d) shows that both sampler has similar performance across varying dimension.
Caution regarding approximation error.
The TV convergence rate in both Theorems 4.5 and 4.7 depends highly on the approximation errors. In our experiments, we find that achieving good approximation error in high-dimensional regime is difficult even for the aforementioned simple low-dimensional Gaussian distribution case. We demonstrate this in a low-dimensional Gaussian example in Figure 6 where we set $k=2$. The figure shows the generated data along the 3rd dimension for both the RF and Stoc-RF samplers. We can see that the growing dimension has an adverse effect on the generation quality. In fact, Stoc-RF appears to be more brittle, presumably due to its inherent stochasticity. The main difficulty in estimation stems from the non-smoothness of $v_{t}(\cdot)$ in (22) along the direction of low-dimensionality when $t\approx 1$. Such non-smoothness is hard to capture via the usual neural network class of functions (e.g., MLP, U-Net) that are typically smooth and necessitate the adoption of more expressive architecture. However, we postpone this direction to future research.
@@
5.2 T2I experiments
We also compare the generation quality of Stoc-RF sampler (21) and the vanilla SDE sampler (20) in Figure 5. The generated plots show that the quality is generally better for Stoc-RF sampler (21) compared to the RF sampler and the vanilla SDE sampler (20).
(a) Prompt: A photo of a cat holding a sign Rectified flow. ($N=50$)
(d) Prompt: A car made of cheese on a wooden table. ($N=80$)
Figure 5: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Columns 1), uniform time-grid (Column 2), Stoc-RF sampler (Column 3), and vanilla SDE sampler (Column 4). Blue color indicates the proposed methods in this paper. The generation quality of Stoc-Rf is generally better out of the four samplers.
6 Conclusions
Appendix A Additional Figures
A.1 Comparison of generation quality of samplers with growing dimension
Figure 6: The generation quality of RF sampler and Stoc-RF sampler along the low-dimensional direction gradually degrades with growing dimension
@@ B.1 Velocity function and its first order identities
With this, Equation (3) can be re-written as
@@
    $\Phi_{t_{i}}(x):=x+\eta_{t_{i}}\widehat{v}_{t_{i}}(x)\quad\text{with $\eta_{t_{i}}=t_{i+1}-t_{i}$},$
@@
Next, we recall that $v_{i}(x)=\frac{x}{t_{i}}+\frac{1-t_{i}}{t_{i}}s_{i}(x)$. Therefore, by (Liu, 2024, Lemma 13) we have
@@
Further, for any vector $v\in\mathbb{R}^{d}$ with $\|v\|_{2}=1$ Assumption 4.4(a) tells us that for $i\leq N-2$
$v^{\top}\frac{\partial\Phi_{i}(x)}{\partial x}v=v^{\top}\left(I+\eta_{i}\frac{\partial\widehat{v}_{i}(x)}{\partial x}\right)v=v^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)v+\eta_{i}v^{\top}\varepsilon_{i}^{J}(x)v$
$\geq v^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)v-\frac{1}{8}\|v\|_{2}^{2}=v^{\top}\left\{\left(1-\frac{\eta_{i}}{1-t_{i}}\right)I+\frac{\eta_{i}t_{i}}{(1-t_{i})^{3}}\operatorname{Cov}_{1\mid t_{i}}(x)\right\}v-\frac{1}{8}\|v\|_{2}^{2}$
$\geq\left(1-h-\frac{1}{8}\right)\|v\|_{2}^{2}>\frac{3}{8},\quad\text{if $h<1/2$.}$
Here, the penultimate relation follows from the positive semi-definiteness of the covariance matrix $\operatorname{Cov}_{1\mid t}(x)$.
@@
Setting $t=\|x\|_{2}$ and $u=x/\|x\|_{2}$, we have $\|\Phi_{i}(x)\|_{2}\geq 3\|x\|_{2}/8-\|\Phi_{i}(0)\|_{2}$. This implies that $\Phi_{i}(\cdot)$ is proper. Then, by Theorem 2.2 of (Ruzhansky and Sugimoto, 2015) shows that $\Phi_{i}(\cdot)$ is a diffeomorphism. Therefore, we can find a unique $x_{i}$ such that ${\Phi}_{i}\left(x_{i}\right)=x_{i+1}$, which in turn allows us to derive
@@
$\mathbb{P}_{\widetilde{\Phi}_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
@@
$=\int_{\widetilde{\Phi}_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\widetilde{\Phi}_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-e^{\xi\left(x_{i},x_{N}\right)}\cdot e^{W\left(x_{i}\right)}\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\widetilde{\Phi}_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{\left(1-e^{\xi\left(x_{i},x_{N}\right)}\right)e^{W\left(x_{i}\right)}+\left(1-e^{W\left(x_{i}\right)}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\int_{x_{i},x_{N}\in\widetilde{\Phi}_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\rm data}}\left\{1-e^{W\left(x_{i}\right)}\right\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\mathrm{d}x_{i}\;{\mathrm{d}x}_{N}$
$\leq\int_{x_{i}\in\widetilde{\Phi}_{i}^{-1}(\mathcal{A})}-W\left(x_{i}\right)p_{X_{i}}\left(x_{i}\right)\mathrm{d}x_{i},$
@@
Lemma B.1 (Lemma 5, (Liang et al., 2025)).
@@
$=\int_{x_{i},x_{N}\in\mathcal{A}\times\mathcal{X}_{\rm data}}\frac{1}{(1-t_{i})^{2}}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{N}\mid X_{i}=x_{i}}(x_{N})p_{X_{i}}(x_{i})\;{\mathrm{d}x}_{N}\;{\mathrm{d}x}_{i}$
@@
Now, recall that $X_{i}\mid X_{N}=x_{N}\sim N(t_{i}x_{N},(1-t_{i})^{2}I)$. Using standard techniques from (Liang et al., 2025, Section B.5), we can obtain bound on (32):
@@
Borrowing standard techniques and results from the proofs of (Liang et al., 2025, Lemma 1) and (Liang et al., 2025, Corollary 1), we have the following results
@@
$\mathbb{P}_{\widetilde{\Phi}_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
$={\mathbb{P}}(\widetilde{\Phi}_{i}(X_{i})\in\mathcal{A},X_{i}\in\mathcal{E}_{i})-{\mathbb{P}}(X_{i+1}\in\mathcal{A})+{\mathbb{P}}(X_{i}\notin\mathcal{E}_{i})$
$\leq{\mathbb{P}}(\widetilde{\Phi}_{i}(X_{i})\in\mathcal{A},\widetilde{\Phi}(X_{i})\in\widetilde{\Phi}(\mathcal{E}_{i}))-{\mathbb{P}}(X_{i+1}\in\mathcal{A})+{\mathbb{P}}(X_{i}\notin\mathcal{E}_{i})$
$\leq{\mathbb{P}}(\widetilde{\Phi}_{i}(X_{i})\in\mathcal{A}\cap\widetilde{\Phi}(\mathcal{E}_{i}))-{\mathbb{P}}(X_{i+1}\in\mathcal{A}\cap\widetilde{\Phi}(\mathcal{E}_{i}))+{\mathbb{P}}(X_{i}\notin\mathcal{E}_{i})$
@@
We also use the Markov’s inequality to control the following:
${\mathbb{P}}(X_{i}\notin\mathcal{E}_{i})={\mathbb{P}}\left\{\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\|\varepsilon_{i}^{\rm J}(x_{i})\|>\frac{1}{8}\right\}\leq\frac{64\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\mathbb{E}\left\{\|\varepsilon_{i}^{\rm J}(x_{i})\|^{2}\right\}\leq\frac{64\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}(\varepsilon^{\rm J,1}_{i})^{2}.$
@@
$\mathbb{P}_{\widetilde{\Phi}_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
@@
Few properties of time scheduling (15):
@@
    For all $i>1$, we have $\frac{\eta_{i}}{1-t_{i}}\leq 2h$ and $\frac{\eta_{i}}{1-t_{i+1}}\leq 2h$.
@@
    For all $i>1$, we have $\frac{\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}=\frac{(1-t_{i})^{2}}{(\frac{1-t_{i}}{\eta_{i}}-1)^{2}}\leq 4h^{2}$, as $\frac{\eta_{i}}{(1-t_{i})}\leq 2h$ and $(1-t_{i})\leq 1$.
@@
    $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}=\sum_{i:0<t_{i+1}\leq 1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}+\sum_{i:t_{i+1}>1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}\leq 4h^{2}N$.
@@
[…]
Now, note that one can estimate $v_{0}(x)=\mu_{p_{X_{1}}}-x$ by $\widehat{v}_{0}(x):=\widehat{\mu}_{1}-x$, where $\widehat{\mu}_{1}=\sum_{i=1}^{n}x_{i,1}/n$ and $n$ is the (training) sample size. In this case, we have
@@
Now, if we substitute all the quantities of the above sampler by its corresponding RF eversions, i.e., we get $\hat{Y}_{\tau}\rightarrow Y_{t}/\sigma_{t},\widehat{s}_{Y^{\prime}_{\tau}}(y)\rightarrow\sigma_{t}\widehat{s}_{t_{i}}(Y_{t_{i}})$, and $\alpha_{\tau}\rightarrow\frac{R_{i}^{2}}{R_{i+1}^{2}}$, then it leads to the following sampler:
@@
We define the auxiliary processes in exact way defined in (Liang et al., 2025, Section C.1).
@@
The DDIM sampler considered in Liang et al. (2025) is the follwing:
@@
Next, using the same set of transformations as in Section B.3 and plugging it in (46), we arrive at the following RF sampler:
@@
Figure 7: Histogram and KDE plot of the transformed time-grid (18) showing U-shaped distribution.
@@
$Y_{t_{i+1}}=\frac{t_{i+1}}{t_{i}}\,Y_{t_{i}}+\frac{(t_{i+1}-t_{i})(1-t_{i})}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})=Y_{t_{i}}+\Delta_{i}\left\{{\frac{Y_{t_{i}}}{t_{i}}+\frac{1-t_{i}}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})}\right\},$ (48)
where $\Delta_{i}=t_{i+1}-t_{i}$. This expression matches exactly the form of the RF sampler in (5). However, the time steps $\Delta_{i}$ are different from those we propose in (15), which we used to obtain the RF convergence guarantees given in Theorem 4.5. Although the time-schedule is still U-shaped as shown in Figure 7.

### v2

@@ 5 Experiments
We perform extensive simulations to corroborate our theoretical findings for both synthetic data and T2I experiments. Codes are available in https://github.com/roysaptaumich/Low-dim-RF.
Figure 3: (a)-(b) Estimated TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF sampler (5) under uniform and non-uniform time-grid under varying $d$. (c)-(d) Estimated TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF and STOC-RF samplers under uniform and non-uniform time-grid under varying $d$.
(a) Prompt: Grape vines in the shape of text ‘open your mind’ sprouting out of a head with flowers and butterflies. DSLR photo. (Better text rendering, $N=50$)
(b) Prompt: A sign that says ”Please refrain from arguing with the chimpanzees”. (Minimal mistakes, $N=50$)
(c) Prompt: Photo of a robot lecturer writing the words ‘Representation Learning’ in cursive on a blackboard, with math formulas and diagrams. (Correct spelling, $N=50$)
Figure 4: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Red frame) and uniform time-grid (Blue frame). Generation quality of RF is better under non-uniform time-grid (15): less hallucinations, enjoys better text rendering and achieves better semantic completeness.
@@
Synthetic data. In this section we demonstrate the usefulness of non-uniform time-grid for generation task from low-dimensional distribution. We set the intrinsic dimension $k=8$, and set our target distribution to be $p_{1}=N(\mu,\Sigma)$, where $\mu=(8,8,\ldots,8)^{\top}\in\mathbb{R}^{d}$ and $\Sigma$ is a block-diagonal matrix with blocks $I_{k}$ and $\boldsymbol{0}_{(d-k)\times(d-k)}$. Tweedie’s formula (Section 2.1) yields
$v_{t_{i}}(x)=\begin{bmatrix}\frac{2t_{i}-1}{\sigma_{t_{i}}^{2}}I_{k}&\boldsymbol{0}_{k\times(d-k)}\\ \boldsymbol{0}_{(d-k)\times k}&-\frac{1}{1-t_{i}}I_{d-k}\end{bmatrix}x.$ (22)
@@
We use (22) for generating new samples from $p_{1}$. We vary $d$ between 10 and 800, and set $N\in\left\{{100,200}\right\}$. We use both a non-uniform time-grid (15) (with $\delta=\min\{1/N,1/d\}$) and a uniform time-grid for the generation task. We generate 2000 samples using (15) and measure the TV distance with respect to another batch of 2000 samples independently generated from the slightly blurred target distribution $\text{Law}((1-\delta)X_{1}+\delta Z)$, where $Z\sim N(0,I_{d})$. We use the discriminative approach proposed in Tao et al. (2024) to estimate the TV distance between the distributions via 10 independent Monte Carlo rounds. Figure 3(a)-(b) show that nonuniform discretization has better adaptation to the low-dimensional structure of $p_{1}$ for large $d$. For the SDE sampler (21), we also compare its TV convergence with that of the deterministic sampler (5) with the non-uniform time grid (15). Figure 3(c)-(d) show that both samplers have similar performance across varying dimensions.
T2I experiments. We perform experiments using Flux (Labs, 2024; Labs et al., 2025) to demonstrate the usefulness of the U-shaped time schedule (15). Figure 4 shows that generation quality of RF is superior under a non-uniform time-grid (15) compared to uniform time-grid. Figure 5 compares the generation quality of STOC-RF sampler (21) and the vanilla SDE sampler (20) and ODE samplers. We see that the generation quality is generally better for STOC-RF. Details are deferred to Appendix A.
(a) Prompt: A photo of a cat holding a sign Rectified flow. ($N=50$)
(d) Prompt: A car made of cheese on a wooden table. ($N=80$)
Figure 5: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Columns 1), uniform time-grid (Column 2), STOC-RF sampler (Column 3), and vanilla SDE sampler (Column 4).
Caution: The TV convergence rates in both Theorem 4.5 and Theorem 4.7 depend highly on the approximation errors. Difficulty may arise in estimating $v_{t}(\cdot)$ along the direction of low-dimensionality when $t\approx 1$ due to non-smoothness (see (22)). Such non-smoothness is usually hard to capture via standard neural networks (e.g., MLP, U-Net) that are typically smooth, and necessitate the adoption of more expressive architecture. We defer this to future research.
6 Conclusions
Acknowledgements
We thank Prof. Eric Vanden-Eijnden at NYU, Prof. Sanjay Shakkottai at UT Austin, and Dr. Christian Rau for helpful discussions and pointing out important references. AR, PS and SR gratefully acknowledge the NSF grants CCF-2019844 and CCF-2505865. PS and SR gratefully acknowledges NSF grant 2217069.
Appendix A Additional Experiments
We also compare the generation quality of STOC-RF sampler (21) and the vanilla SDE sampler (20) in Figure LABEL:fig:_Flux_SDE/RF_experiments. The generated plots show that the quality is generally better for STOC-RF sampler (21) compared to the RF sampler and the vanilla SDE sampler (20).
@@ B.1 Velocity function and its first order identities
With this, true drift be re-written as
@@
    $\Phi_{t_{i}}(x):=x+\eta_{i}\widehat{v}_{t_{i}}(x)\quad\text{with $\eta_{i}=t_{i+1}-t_{i}$},$
@@
Next, we recall that $v_{i}(x)=\frac{x}{t_{i}}+\frac{1-t_{i}}{t_{i}}s_{i}(x)$. Therefore, by Liu (2024, Lemma 13) we have
@@
Further, for any vector $u\in\mathbb{R}^{d}$ with $\|u\|_{2}=1$ Assumption 4.4(a) tells us that for $i\leq N-2$
$u^{\top}\frac{\partial\Phi_{i}(x)}{\partial x}u=u^{\top}\left(I+\eta_{i}\frac{\partial\widehat{v}_{i}(x)}{\partial x}\right)u=u^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)u+\eta_{i}u^{\top}\varepsilon_{i}^{J}(x)u$
$\geq u^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)u-\frac{1}{8}\|u\|_{2}^{2}=u^{\top}\left\{\left(1-\frac{\eta_{i}}{1-t_{i}}\right)I+\frac{\eta_{i}t_{i}}{(1-t_{i})^{3}}\operatorname{Cov}_{1\mid t_{i}}(x)\right\}u-\frac{1}{8}\|u\|_{2}^{2}$
$\geq\left(1-h-\frac{1}{8}\right)\|u\|_{2}^{2}>\frac{3}{8},\quad\text{if $h<1/2$.}$
Here, the penultimate relation follows from the positive semi-definiteness of the covariance matrix $\operatorname{Cov}_{1\mid t}(x)$ and the fact that $\frac{\eta_{i}}{1-t_{i}}=\frac{ht_{i}}{1-t_{i}}{\mathbbm{1}}\left\{{t_{i}\leq 1/2}\right\}+\frac{h(1-t_{i+1})}{1-t_{i}}{\mathbbm{1}}\left\{{t_{i}>1/2}\right\}\leq h$ (see Lemma D.2). Here $\mathbbm{1}$$\left\{{\cdot}\right\}$ denotes the indicator function.
@@
Setting $t=\|x\|_{2}$ and $u=x/\|x\|_{2}$, we have $\|\Phi_{i}(x)\|_{2}\geq 3\|x\|_{2}/8-\|\Phi_{i}(0)\|_{2}$. This implies that $\Phi_{i}(\cdot)$ is proper. Then, by Theorem 2.2 of Ruzhansky and Sugimoto (2015) shows that $\Phi_{i}(\cdot)$ is a diffeomorphism. Therefore, we can find a unique $x_{i}$ such that ${\Phi}_{i}\left(x_{i}\right)=x_{i+1}$, which in turn allows us to derive
@@
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
@@
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-e^{\xi\left(x_{i},x_{N}\right)}\cdot e^{W\left(x_{i}\right)}\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{\left(1-e^{\xi\left(x_{i},x_{N}\right)}\right)e^{W\left(x_{i}\right)}+\left(1-e^{W\left(x_{i}\right)}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\int_{x_{i},x_{N}\in\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\rm data}}\left\{1-e^{W\left(x_{i}\right)}\right\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\mathrm{d}x_{i}\;{\mathrm{d}x}_{N}$
$\leq\int_{x_{i}\in\Phi_{i}^{-1}(\mathcal{A})}-W\left(x_{i}\right)p_{X_{i}}\left(x_{i}\right)\mathrm{d}x_{i},$
@@
Lemma B.1 (Lemma 5, Liang et al. (2025)).
@@
Now, recall that $X_{i}\mid X_{N}=x_{N}\sim N(t_{i}x_{N},(1-t_{i})^{2}I)$. Using standard techniques from Liang et al. (2025, Section B.5), we can obtain bound on (32):
@@
Borrowing standard techniques and results from the proofs of Liang et al. (2025, Lemma 1) and Liang et al. (2025, Corollary 1), we have the following results
@@
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
@@
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
@@
Recalling properties of time scheduling (15) (see Lemma D.2):
@@
    For all $i>1$, we have $\frac{\eta_{i}}{1-t_{i}}\leq h$ and $\frac{\eta_{i}}{1-t_{i+1}}\leq h$.
@@
    For all $i>1$, we have $\frac{\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}=\frac{(1-t_{i})^{2}}{(\frac{1-t_{i}}{\eta_{i}}-1)^{2}}\leq h^{2}$, as $\frac{\eta_{i}}{(1-t_{i})}\leq h$ and $(1-t_{i})\leq 1$.
@@
    $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}=\sum_{i:0<t_{i}<1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}+\sum_{i:t_{i}\geq 1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}\leq 4h^{2}N$.
@@
$\quad+A_{1}\delta^{2}{\operatorname{Tr}(\text{Cov}_{1})}+A_{2}\delta^{2}(\varepsilon^{\rm v}_{0})^{2}+A_{3}\delta^{2}(\varepsilon^{\rm J,1}_{0})^{2}+A_{4}(\varepsilon^{\rm v}_{0}+\varepsilon^{\rm J,1}_{0}+\varepsilon^{\rm J,2}_{0}+\varepsilon^{\rm H}_{0}),$
where $\text{Cov}_{1}:=\text{Cov}_{X_{1}\sim p_{1}}(X_{1})$. Now, note that one can estimate $v_{0}(x)=\mu_{p_{X_{1}}}-x$ by $\widehat{v}_{0}(x):=\widehat{\mu}_{1}-x$, where $\widehat{\mu}_{1}=\sum_{i=1}^{n}x_{i,1}/n$ and $n$ is the (training) sample size. In this case, we have
@@
Now, if we substitute all the quantities of the above sampler by their corresponding RF versions, i.e., we get $\hat{Y}_{\tau}\rightarrow Y_{t}/\sigma_{t},\widehat{s}_{Y^{\prime}_{\tau}}(y)\rightarrow\sigma_{t}\widehat{s}_{t_{i}}(Y_{t_{i}})$, and $\alpha_{\tau}\rightarrow\frac{R_{i}^{2}}{R_{i+1}^{2}}$, then it leads to the following sampler:
@@
We define the auxiliary processes in exact way defined in Liang et al. (2025, Section C.1).
@@
    Note that, the above update rule is just the analog of $\{\hat{Y}_{\tau}\}_{\tau}$ process with the true score.
@@
The DDIM sampler considered in Liang et al. (2025) takes the form
@@
Using the same transformations detailed in Section B.3 and plugging them in (46), we arrive at the RF sampler
@@
(a) Histogram and kernel density estimation (KDE) plot of time-grid (15) showing U-shaped distribution.
(b) Histogram and KDE plot of the transformed time-grid (18) showing U-shaped distribution.
Figure 7: Histogram of non-uniform time grids.
@@
$Y_{t_{i+1}}=\frac{t_{i+1}}{t_{i}}\,Y_{t_{i}}+\frac{(t_{i+1}-t_{i})(1-t_{i})}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})=Y_{t_{i}}+\Delta_{i}\underbrace{\left\{{\frac{Y_{t_{i}}}{t_{i}}+\frac{1-t_{i}}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})}\right\}}_{=\widehat{v}_{t_{i}}(Y_{t_{i}})},$ (48)
where $\Delta_{i}=t_{i+1}-t_{i}$. This expression matches exactly the form of the RF sampler in (5). However, the time steps $\Delta_{i}$ are different from those we propose in (15), which we used to obtain the RF convergence guarantees given in Theorem 4.5. Although the time-schedule is still U-shaped as shown in Figure 7 (b), however, it is not symmetric and, as a result, might take more discretization steps at either the beginning or the end of the sampling stage. Whether a symmetric or asymmetric schedule is better remains an open problem. Furthermore, we observe that empirically it appears to be relatively sensitive to the choice of the parameters $c_{0}$ and $c_{1}$ in (18).